OpenStudy (anonymous):

DOUBLE INTEGRAL: the integral from x=0 to x=1 of the integral y=0 to y=2 of the function xy(e^((x^2)y))

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \int\limits_{0}^{2}xye ^{x ^{2} y}dydx\]

OpenStudy (anonymous):

Do first integral\[x \int\limits_{0}^{2}ye ^{x2y}dy\]

OpenStudy (anonymous):

Actually, shuffle it around do x first\[y \int\limits_{0}^{1}xe ^{x ^{2}y}dx\]Do this first you can use u sub.

OpenStudy (anonymous):

wow, thanks a lot =)